​Order: Phenytoin (Dilantin) 15​ mg/kg loading dose IV push stat. Phenytoin is supplied 50​ mg/mL and is to be administered at 25​ mg/min. How many mL would you administer over how many minutes if the client weighs 65​ kg?

A) 30 mL to be given over 30 minutes

B) 19.5 mL over 20 minutes

C) 30 mL to be given over 1 hour

D) 19.5 mL over 39 minutes
Answer #1

Dose = 15 mg/kg

For 65 kg, 15*65 = 975 mg has to be given.

Max rate is 25mg/min

So for giving 975 mg, 975/25 min is required = 39 minutes.

50 mg is present in 1 mL.

975 mg is present in 975/50 mL = 19.5 mL.

So option D is correct